OpenStudy (anonymous):

The perimeter of a rectangle is 32 feet. Find the dimensions of the rectangle if the length is 4 feet longer than three times the width. Then find the area of the rectangle.

OpenStudy (anonymous):

2(x+y)=32 x+4=y

OpenStudy (anonymous):

Could you explain? Im confused :P

OpenStudy (anonymous):

where x and y are the length and width

OpenStudy (anonymous):

use substitution to solve for x and y

OpenStudy (anonymous):

So would I do 3x+4=32? 4 more than 3 times the width?

OpenStudy (anonymous):

Sorry, I really dont get this stuff.

OpenStudy (anonymous):

Oh wait, I get it, nevermind, thank you!

OpenStudy (anonymous):

put x+4=y into 2(x+y)=32 to get 2(x+x+4)=32

OpenStudy (anonymous):

2(2x+4)=32

OpenStudy (anonymous):

after you solve it, substitute x back into one of the equations to solve for y